# Hierarchical Vision-Language Prompting for Long-form Video Understanding

### v2

@@ 4.1 Experimental Setup
Datasets: Given the training free nature of our framework, we do not utilize any video datasets for training. Datasets are used purely for evaluation. We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [Mangalam2023EgoSchemaAD], NExT-QA [dataset_xiao2021nextqa], and IntentQA [li2023intentqa_cavir]. EgoSchema is a very long-form video question-answering dataset derived from Ego4D. For each question, EgoSchema requires the correct answer to be selected between five given options based on a three-minute-long video clip. We evaluate on both its 500 subset and the 5000 fullset.
@@
NExT-QA is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer. We evaluate on its validation set containing 4996 questions over 570 videos. The third dataset, IntentQA [li2023intentqa_cavir] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions. It consists of 16k multiple-choice questions based on 4.3k videos, each with 5 answer choices. These questions are further classified as why?, how? or before/after, based on the appearance of the same keywords in the question.
VLM and LLM Model Choices: In line with prior state-of-the-art work [wang2024lifelongmemory, zhang2023llovi, wang2023vamos] we utilize GPT-4 API for the VLM and LLM in the latter two stages of our framework. This allows fairer comparison against these methods. In the case of the HKS module, we use the same VLM / LLM for keywords used in CKD and VLM used in FKD.
4.2 Quantitative and Qualitative Evaluation
Table 1: Results on EgoSchema [Mangalam2023EgoSchemaAD]: We present comparisons with state-of-the-art models on EgoSchema subset (500-videos) and fullset (5000-videos). LVNet shows a strong zero-shot performance using only a fraction of frames compared to most methods.
Model Captions Subset (%) Fullset (%)
finetuned
ImageViT [papalampidi2023simple] 16 40.8 30.9
ShortViViT [papalampidi2023simple] 32 47.9 31.0
MC-ViT-L [balavzevic2024mcvit]   128+ 62.6 44.4
LongViViT [papalampidi2023simple] 256 56.8 33.3
zero-shot
VIOLET[fu2023violet] 5 - 19.9
mPLUG-Owl [ye2023mplug] 5 - 31.1
VideoAgent [Wang2024VideoAgentLV] 8.4 60.2 54.1
MVU [rana2024mvu] 16 60.3 37.6
SeViLA [yu2024sevila] 32 25.7 22.7
LifelongMemory [wang2024lifelongmemory] 39 26.4 -
InternVideo [wang2022internvideo] 90 - 32.1
FrozenBiLM [yang2022frozenblim] 90 - 26.9
LifelongMemory [wang2024lifelongmemory] 90 68.0 62.1
LLoVi [zhang2023llovi] 180 57.6 50.3
LangRepo [Kahatapitiya2024langrepo] 180 66.2 41.2
Vamos [wang2023vamos] - - 48.3
LVNet (ours) 12 66.0 61.1
We first evaluate on the EgoSchema dataset and present our results in Table 1. Our framework outperforms most prior work to create a stronger baseline with a few frame captions. While LifelongMemory [wang2024lifelongmemory] achieves the absolute best zero-shot performance on the subset at 90 captions (68.0%), its performance drops significantly at 39 captions (26.4%). In contrast, LVNet shows a 66.0% accuracy, merely at 12 captions, even outperforming VideoAgent [Wang2024VideoAgentLV] (60.2% at 8.4 captions) and MVU [rana2024mvu] (60.3% at 16 captions). We achieve 61.1% for the fullest, the highest among the models utilizing approximately 12 frames. Here, we generate captions independently without conditioning on previous captions and present them to the LLM in a system-user-prompt style.
Table 2: Results on NExT-QA [xiao2021nextqa]: We compare LVNet against prior methods on NExT-QA validation set, highlighting standard splits: causal, temporal and descriptive. Our method shows state-of-the-art zero-shot performance, even competing with finetuned models.
Model Captions  Causal (%)  Temporal (%)  Descriptive (%)  All (%)
finetuned
HiTeA [ye2023hitea] 8 62.4 58.3 75.6 63.1
SeViTFiD [kim2023sevit] 10 - - - 60.6
BLIP-2 [li2023blip2] 10 70.1 65.2 80.1 70.1
LLama-VQA [ko2023llama-vqa] 10 72.7 69.2 75.8 72.0
CoVGT [xiao2023covgt] 32 58.8 57.4 69.3 60.0
SeViLA [yu2024sevila] 32 74.2 69.4 81.3 73.8
InternVideo [wang2022internvideo] 90 62.5 58.5 75.8 63.2
MC-ViT-L [balavzevic2024mcvit] 128+ - - - 65.0
Vamos [wang2023vamos] - 72.6 69.6 78.0 72.5
zero-shot
VideoAgent [Wang2024VideoAgentLV] 8.2 72.7 64.5 81.1 71.3
MVU [rana2024mvu] 16 55.4 48.1 64.1 55.2
VFC [momeni2023vfc] 32 45.4 51.6 64.1 51.5
SeViLA [yu2024sevila] 32 61.3 61.5 75.6 63.6
InternVideo [wang2022internvideo] 90 43.4 48.0 65.1 49.1
LLoVi [zhang2023llovi] 90 69.5 61.0 75.6 67.7
LangRepo [Kahatapitiya2024langrepo] 90 64.4 51.4 69.1 60.9
ViperGPT [suris2023vipergpt] - - - - 60.0
LVNet (ours) 12 75.0 65.5 81.5 72.9
@@ 4.2 Quantitative and Qualitative Evaluation
We next evaluate on the Next-QA dataset and present results in Table 2. This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs. Our approach achives state-of-the-art performance on this benchmark outperforming prior work by a considerable margin. In fact, our LVNet with just 12 frame captions achieves 72.9% overall accuracy in zero-shot, outperforming VideoAgent [Wang2024VideoAgentLV] (71.3% at 8.2 captions) and LLoVi [zhang2023llovi] (67.7% at 90 captions). It is also extremely competitive with finetuned models such as LLama-VQA [ko2023llama-vqa] (72.0% at 10 captions), SeViLA [yu2024sevila] (73.8% at 32 captions) and Vamos [wang2023vamos] (72.5%), while having no video level training.
@@
Table 3: Results on IntentQA [li2023intentqa_cavir]: We compare LVNet against prior methods on IntentQA test set, highlighting standard splits: why?, how? and before/after. Our method shows state-of-the-art performance in zero-shot, with a significant margin.
Model Captions Why? (%) How? (%) Before/After (%) All (%)
finetuned
HQGA [xiao2022hqga] 16 48.2 54.3 41.7 47.7
VGT [xiao2022vgt] 32 51.4 56.0 47.6 51.3
BlindGPT [ouyang2022blindgpt] - 52.2 61.3 43.4 51.6
CaVIR [li2023intentqa_cavir] - 58.4 65.5 50.5 57.6
Vamos [wang2023vamos] - 69.5 70.2 65.0 68.5
zero-shot
SeViLA [yu2024sevila] 32 - - - 60.9
LLoVi [zhang2023llovi] 90 68.4 67.4 51.1 64.0
LangRepo [Kahatapitiya2024langrepo] 90 62.8 62.4 47.8 59.1
LVNet (ours) 12 75.2 71.6 60.8 71.1
We finally evaluate on IntentQA [li2023intentqa_cavir] dataset and report these results in Table 3. Our approach achieves state-of-the-art results on this benchmark as well outperforming all prior work, both zero-shot and finetuned. In fact, LVNet achieves a significant boost (+6.5%) over the closest zero-shot baseline LLoVi [zhang2023llovi] (64.0% at 90 captions), showing a 70.5% accuracy with just 12 frame captions. It also extends the state-of-the-art finetuned performance by +2.0%, compared to Vamos [wang2023vamos] (68.5%).
Figure 3: Qualitative example: We illustrate a challenging long-video QA scenario from EgoSchema [Mangalam2023EgoSchemaAD]. We consider an input of 900 frames, which first get clustered into scenes and subsampled to retain around 280 frames. Next, the Coarse Keyframe Detector selects only 32 frames out of them, based on the alignment with keywords (Here, keywords are extracted based on answer options, via an LLM). Such coarse keyframes are then ordered according to confidence values, and grouped into 4 groups of 8 frames each. These groups go through visal templating (i.e. simple concatenation across space) and fed into a VLM for Fine Keyframe Detection, resulting in just 12 frames.
@@ 4.3 Ablations
Table 4: Ablations on EgoSchema [Mangalam2023EgoSchemaAD]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot video VQA.
Keyframe sampler Seq. captions #captions Acc. (%)
Uniform Sampling (LLoVi) 180 57.6
HKS (w/o SVL) 12 63.0
HKS 12 63.4
@@
Temporal order 62.9
Confidence order 63.4
@@
GPT-3.5 59.2
GPT-4 61.1
GPT-4o 63.4
Frames 6.4 8.4 9 11 12 16
VideoAgent 58.4 60.2 - 57.4 - -
LVNet (ours) - - 64.6 - 66.0 65.8
In this section, we present ablations to key design decisions such as #frames, #words, prompting, and hierarchical processing. The results for these ablations are presented in Table 4. In all ablations, we consider a subset of 150 videos from EgoSchema [Mangalam2023EgoSchemaAD] for evaluation.
Keyframe sampler:
We evaluate our Hierarchical Keyframe Sampling strategy (HKS) against the established uniform sampling baseline from [zhang2023llovi]. Our results clearly indicate the importance of our HKS module.
Sequential captioning:
In our SVL module, we consider a sequential captioner, that considers previous captions when captioning the current frame. It outperforms the common parallel frame-wise captioning by $+0.4\%$ showing that the SVL can generate better captions, possibly reducing redundant information among captions and enabling a form of temporal modeling.
Visual Templating order:
In Visual Templating, grouping the frames based on confidence is better than preserving temporal order. It can allow sampling of challenging keyword frames, instead of only selecting high confidence keyframes.
@@
Here, we see that more-recent GPT-4o outperforms both GPT-4 and GPT-3.5. Since GPT-4o is also cheaper compared to GPT-4, we consider it as our default choice.
Number of frame captions:
We ablate different number of frame captions used in our setup compared to VideoAgent [Wang2024VideoAgentLV]. Our method shows more-stable performance.
5 Conclusion
We proposed an approach for Long-form Video Question Answering (LVQA) that achieves state-of-the-art results across 3 benchmarks datasets. Our propsoed Hierarchical Keyframe Selector and Sequential Visual LLM modules demonstrate the usefulness of keyframe selection and sequence-aware frame captioning. We also highlight the zero-shot capability for long-form video understanding of our LVNet framework requiring zero video level training, and our experiments demonstrate its clear advantage over prior work.

### v3

@@ 4.1 Experimental Setup
Datasets: Given the training free nature of our framework, we do not utilize any video datasets for training. Datasets are used purely for evaluation. We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [24], IntentQA [22], and NExT-QA [41]. The first dataset, EgoSchema, is a very long-form video question-answering dataset with 5031 questions. For each question, EgoSchema requires the correct answer to be selected between five given options based on a three-minute-long video clip. The second dataset, IntentQA [22] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions. It consists of 16k multiple-choice questions which are classified as Why?, How? or Before/After (B./A.). The third dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer. These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos.
Model Choices: For the HKS module, we use the ResNet-18 [12] for the TSC, CLIP-B/16 [29] for the CKD and GPT-4o for the FKD. We select ResNet-18 and CLIP-B/16 due to their smaller models sizes—0.01B and 0.12B, respectively–which are significantly lighter compared to GPT-4o, whose model size is expected to be on the scale of 100B-1T. This makes them well-suited for filtering dense frames efficiently. In line with previous state-of-the-art work [39, 52, 36], we employ GPT API, especially GPT-4o, for both VLM and LLM. This choice is driven by its cost-effectiveness and lighter computational requirements compared to GPT-4. GPT-4o is used as the VLM for generating captions and as the LLM for answering questions in our framework.
Table 1: Combined Results on IntentQA [22], EgoSchema [24], and NExT-QA [41]. We compare LVNet against prior zero-shot models across three datasets, highlighting different task splits. The models are ordered based on the number of captions processed before answering the question. LVNet achieves state-of-the-art accuracies of 71.7%, 61.1%, and 72.9% on the three datasets, respectively, using just 12 frames compared to the models using the similar number of captions. Models with video-caption pretraining or utilizing significantly more captions than the 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency.
Model IntentQA EgoSchema NExT-QA
Cap. Why? (%) How? (%) B./A. (%) All (%) Cap. Acc. (%) Cap. Cau. (%) Tem. (%) Des. (%) All (%)
Vamos [36] - - - - - - 48.3 - - - - -
IG-VLM [17] - - - - 65.3 - 59.8 - 69.8 63.6 74.7 68.6
VideoLLaMA 2 [5] - - - - - - 53.3 - - - - -
InternVideo2 [38] - - - - - - 60.2 - - - - -
Tarsier [35] - - - - - - 61.7 - - - - 79.2
VIOLET [10] - - - - - 5 19.9 - - - - -
mPLUG-Owl [46] - - - - - 5 31.1 - - - - -
VideoAgent [37] - - - - - 8.4 54.1 8.2 72.7 64.5 81.1 71.3
MVU [30] - - - - - 16 37.6 16 55.4 48.1 64.1 55.2
MoReVQA [25] - - - - - 30 51.7 16 70.2 64.6 - 69.2
TraveLER [37] - - - - - - - $\sim$25 70.0 60.5 78.2 68.2
VFC [26] - - - - - - - 32 45.4 51.6 64.1 51.5
SeViLA [48] 32 - - - 60.9 32 22.7 32 61.3 61.5 75.6 63.6
ProViQ [6] - - - - - 60 57.1 60 - - - 64.6
VideoTree [40] $\leq$56 - - - 66.9 63.2 61.1 $\leq$56 75.2 67.0 81.3 73.5
FrozenBiLM [45] - - - - - 90 26.9 - - - - -
LifelongMemory [39] - - - - - 90 62.1 - - - - -
LangRepo [16] 90 62.8 62.4 47.8 59.1 180 41.2 90 64.4 51.4 69.1 60.9
LLoVi [52] 90 68.4 67.4 51.1 64.0 180 50.3 90 69.5 61.0 75.6 67.7
LVNet (ours) 12 75.0 74.4 62.1 71.7 12 61.1 12 75.0 65.5 81.5 72.9
4.2 Evaluation
Quantitative Results:
We evaluate LVNet on the IntentQA, EgoSchema, and NExT-QA dataset and present our results in Table 1. Models with video-caption pretraining are de-emphasized in grey to ensure fairness with image-level pertaining. Models utilizing significantly more captions than the 12 frames are downplayed in light green to consider caption efficiency. For EgoSchema, we achieve 61.1% on the fullest, the highest among the models utilizing approximately 12 captions. This result outperforms VideoAgent, the next best model using 8.4 captions, by 7%. To ensure a fair comparison, we ablate the number of captions for LVNet in LABEL:tab:ablation:numberofframecaptions, observing that LVNet with 8 captions still outperforms VideoAgent with 8.4 captions by 4.2% in the EgoSchema subset. The superior accuracy of LVNet over other keyframe selectors, such as those used in VideoAgent and TraveLER, is discussed in Section 3.4.
We next evaluate on the IntentQA[22] dataset. Our approach achieves state-of-the-art results on this benchmark, outperforming all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames. In fact, LVNet shows a substantial improvement (+4.8%) over the next best model, VideoTree [40] (66.9% with 90 captions), achieving 71.7% accuracy with only 12 frame captions.
@@ 4.2 Evaluation
We finally evaluate on the NExT-QA[41] dataset. This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs. Our approach achives state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions. In fact, our LVNet with just 12 frame captions achieves 72.9% overall accuracy, outperforming VideoAgent [37] (71.3% at 8.2 captions) by 1.6%.
@@
Qualitative Analysis of the Hierarhical Keyframe Selector:
We compare the open-ended responses of LVNet and the uniform sampling method in Figure 4 to understand the effectiveness of the hierarchical keyframe selector in LVNet. The frames chosen by LVNet and the naive uniform sampling method are indicated by blue and red checkmarks in the images, respectively. LVNet selects frames at 5, 69, and 135 seconds by executing the hierarchical keyframe selector and generates captions based on those frames. When we feed the concatenated captions to the LLM to answer the given question: "Based on the video, what are the three main types of tools that C uses…" in an open-ended manner, the output identifies two main activities: welding torches and measuring tapes, among the three main activities described in Option 3 (welding handle, hammer, and measuring tape), which is the correct answer. This leads LVNet to choose the correct option.
In contrast, the uniform sampling method selects frames at 0, 16, and 32 seconds and generates captions based on those frames. Similarly, when we feed the concatenated captions to the LLM to answer the same question, the output identifies only one activity—welding tools—resulting in the selection of the incorrect option. This example highlights the importance of keyframe selection and demonstrates the effectiveness of hierarchical keyframe selection in LVNet.
Figure 4: Open-ended Responses from LVNet vs Uniform Sampling: The frames chosen by LVNet and the naive uniform sampling method are indicated with blue and red checkmarks, respectively. LVNet identifies both welding torches and measuring tapes, choosing the correct option, whereas uniform sampling only detects welding tools and selects the incorrect answer. The blue, red, and purple highlights correspond to the three main activities in the video—welding a handle, using a hammer, and using a measuring tape, respectively.
@@ 4.3 Ablations
Table 2: Ablations on EgoSchema [24]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot video VQA.
@@
Temporal 65.2
Hybrid (Confidence + Temporal) 68.2
TSC CKD FKD Acc.
68.2
65.8
64.5
62.6
Frames 6.4 8 8.4 11 12 16
VideoAgent[37] 58.4 - 60.2 57.4 - -
LVNet (ours) - 64.4 - - 68.2 67.8
In this section, we present ablations on key design decisions such as the sorting order in FKD, the number of frames for captions, and the effect of different components in HKS. In all ablations, we use a subset of EgoSchema [24], composed of 500 videos. Additional ablations about Choice of LLM and Effect of Patch Size on Keyword Matching in CKD are in Appendix A.1
Visual Templating Order:
In visual templating, prioritizing frames by keyword confidence scores followed by reordering low-confidence frames based on timestamp is more effective than relying solely on temporal order. In this hybrid approach, high-confidence frames are selected from shorter segments by sampling three keyframes per set of eight, grouped by confidence rather than timespan. Conversely, low-confidence keyframes, which are crucial but more visually challenging for keyword matching, are sampled from broader video sections, with three frames per set of eight grouped by timestamp. This hybrid approach outperforms purely temporal ordering by 3%.
Number of Frame Captions:
We conducted an ablation study on the number of frame captions used in our setup, comparing it to VideoAgent [37], which operates with a similarly small number of captions. Our findings show that LVNet achives its highest accuracy 68.2% with 12 captions, while even its lowest accuracy 64.4% with 8 captions still surpasses VideoAgent’s best accuracy 60.2% with 8.4 captions.
Effect of Hierarchical Keyframe Modules:
This table demonstrates the impact of incrementally adding the temporal scene clustering (TSC), coarse keyframe detector (CKD), and fine keyframe detector (FKD) modules. Without any of these modules, the model relies on uniform sampling and achieves 62.6%. When TSC is added and 12 frames are selected uniformly, the accuracy increases to 64.5%. Adding both TSC and CKD raises the accuracy to 65.8%. Finally, incorporating all three modules—TSC, CKD, and FKD—into the model, which is LVNet, results in an accuracy of 68.2%. This demonstrates the importance of including all modules in LVNet for optimal performance.
5 Conclusion
We proposed a novel approach for Long-form Video Question Answering (LVQA) that achieves state-of-the-art performance compared to the model using the similar-scale captions across 3 benchmarks datasets. Our Hierarchical Keyframe Selector demonstrates the effectiveness of keyframe selection in understanding a very long-form video QA. Additionally, we highlight the zero-shot capability for long-form video comprehension of our LVNet framework, which requires no video-level training. Our experiments showcase its significant advantage over previous methods.
Appendix
Appendix A.1 Additional Ablations
In this section, we present additional experiments conducted to inform the LVNet’s design. We have tested different LLMs and experimented with various scales of the visual feature map.
Table A.3: Ablations on EgoSchema [24]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA.
@@ Appendix A.1 Additional Ablations
GPT-3.5 61.0
GPT-4 65.4
GPT-4o 68.2
Patch Size Acc. (%)
1x1 63.6
7x7 66.2
14x14 68.2
@@
Our results show that GPT-4o outperforms both GPT-4 and GPT-3.5 by 2.8% and 7.2%, respectively. Given that GPT-4o is more cost-effective and lightweight compared to GPT-4, we have selected it as our default LLM.
Effect of Patch Size on Keyword Matching in CKD:
Since keywords can represent activities spanning the entire image or confined to a small region, we adjust the resolution of the visual feature map output from the spatially aware contrastive image pre-training (CLIP) network [29] to match keywords. Our findings show that higher resolutions lead to better accuracy. In LVNet, we use a 14$\times$14 feature map and determine the confidence level of the keyword by selecting the maximum value between the 14$\times$14 patches and the keyword’s text embedding.
Appendix A.2 Algorithms in Detail
Our algorithms are presented in full detail in Algorithm 1, Algorithm 2, and Algorithm 3. TSC in Algorithm 1 extracts per-frame visual features using ResNet-18, followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets. Within each of the $n$ sets, we uniformly sample $\leq\tau$ frames, obtaining a total of $T_{a}\leq\tau\times n$ frames. For example, LVNet sets $\psi=5,\lambda=12,\tau=18$, resulting in approximately $n\sim 25$ and $T_{a}\sim 390$ on the EgoSchema dataset. CKD in Algorithm 2 selects top $L$ frames based on similarity/confidence scores, which are calculated using cosine similarity between frames and keywords with CLIP-B/16. LVNet employs $L=32,len(K)\leq 25$ on the EgoSchema dataset. FKD in Algorithm 3 sorts frames and their corresponding keywords by confidence scores, and reorder the $K$ frames with the lowest scores temporally. It groups frames sequentially into visual templates, each consisting of $N$ frames. From each template, the $M$ frames and keywords most relevant among the $N$ pairs are selected using GPT-4o. We set $L=32,K=16,N=8,M=3$.
1:Require: ResNet-18 [11] pretrained on imagenet dataset $f$, frame list $\textbf{List}_{frame}$, image index list $\textbf{List}_{index}\in\{1,\dots,N\}$, minimum number of list length $\psi$, temperature $\lambda$, number of sample $\tau$, function to find index of x in list w index(x, w), and function to sort list sort(List)
2:for all $img^{i}$ in $\textbf{List}_{frame}$ do
3:   $\mathbf{F}^{i}\leftarrow f(img^{i})$
4:   $\textbf{List}_{feat}$.insert($\mathbf{F}^{i}$)
5:end for
6:for all $\mathbf{F}^{i}$ in $\textbf{List}_{feat}$ do
7:   $\textbf{List}_{dist}\leftarrow\frac{\sum_{y}\sum_{x}{\sqrt{(\mathbf{F_{i}}-\textbf{List}_{feat})^{2}}}}{x\times y}$
8:   $\textbf{M}_{dist}$.insert($\textbf{List}_{dist}$)
9:end for
10:while length of $\textbf{List}_{index}>\psi$ do
11:   $\textbf{List}_{sample}\leftarrow\emptyset$
12:   $\textbf{List}_{\delta}\leftarrow\emptyset$
13:   $i\leftarrow\textbf{List}_{index}$.pop(00)
14:   $\textbf{p}^{i}\leftarrow\texttt{softmax}(\textbf{M}^{i}_{dist})$
15:   $\mu_{\mathbf{\textbf{p}^{i}}},\sigma_{\mathbf{\textbf{p}^{i}}}\leftarrow\texttt{mean}(\textbf{p}^{i}),\texttt{std}(\textbf{p}^{i})$
16:   $\beta\leftarrow\mu_{\mathbf{\textbf{p}^{i}}}-\sigma_{\mathbf{\textbf{p}^{i}}}\sum_{i=0}{e^{1-i/\lambda}}$
17:   for all prob in $\textbf{p}^{i}$ do
18:      if prob < $\beta$ then
19:         $\textbf{List}_{selected}$.insert($\textbf{index}(\texttt{prob},\textbf{p}^{i})$)
20:      end if
21:   end for
22:   for all $\gamma$ in $\textbf{List}_{selected}$ do
23:      $\delta$leftarrow $\gamma~{}th$ value in $\textbf{List}_{index}$
24:      $\textbf{List}_{\delta}$.insert($\delta$)
25:      $\textbf{List}_{index}$.pop($\gamma$)
26:   end for
27:   $\textbf{List}_{\delta}$.insert($i$)
28:   $\textbf{List}_{sample}$leftarrow sample $\tau$ items from $\textbf{List}_{\delta}$
29:   $\textbf{sort}(\textbf{List}_{sample})$
30:   for all $frame^{j}$ in $\textbf{List}_{frame}$ do
31:      if $j$ in $\textbf{List}_{sample}$ then
32:         Outputs.insert($frame^{j}$)
33:      end if
34:   end for
35:end while
Algorithm 1 Temporal Scene Clustering
1:Require: keyword set $\mathbf{K}$, image set $\mathbf{I}$, total length of selected image set $L$, function to calculate similarity matrix $\textbf{sim}(\mathbf{K},\mathbf{I})$, function to sort similarity matrix and return indices $\textbf{sort}(\mathbf{S})$
2:$\mathbf{S}\leftarrow\textbf{sim}(\mathbf{K},\mathbf{I})$
3:$\mathbf{S}_{sorted},\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
4:Initialize $\mathbf{P}_{best}$ as an empty list
5:Initialize $\mathbf{I}_{selected}$ as an empty set
6:while length of $\mathbf{I}_{selected}<L$ do
7:   for $k\in\mathbf{K}$ do
8:      for $i\in\mathbf{I}$ do
9:         $i_{index}\leftarrow\mathbf{idx}_{sorted}[k][i]$
10:         if $i_{index}$ not in $\mathbf{I}_{selected}$ then
11:            $\mathbf{P}_{best}$.insert($k,i_{index}$)
12:            $\mathbf{I}_{selected}$.insert($i_{index}$)
13:            break
14:         end if
15:      end for
16:      if length of $\mathbf{I}_{selected}\geq L$ then
17:         break
18:      end if
19:   end for
20:end while
21:return $\mathbf{P}_{best}$
Algorithm 2 Keyword-Image Matching Process in CKD
[…]
2:$\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
[…]
4:$\mathbf{idx}_{temporal}\leftarrow\textbf{temporal\_order}(\mathbf{idx}_{low\_sim})$
[…]
6:$\mathbf{I}_{ordered},\mathbf{K}_{ordered}\leftarrow\mathbf{I}[\mathbf{idx}_{final}],\mathbf{K}[\mathbf{idx}_{final}]$
[…]
8:for each $\mathbf{set}\in\mathbf{sets}$ do
9:   $\mathbf{I}_{selected}\leftarrow\texttt{select\_top\_M}(\mathbf{set},M)$
10:end for
11:return $\mathbf{I}_{selected}$
Algorithm 3 Fine Keyframe Detection Process (FKD)
Appendix A.3 Prompting: Fine Keyframe Detector
We prompt the VLM to select frames that are most compatible with the list of given keywords. Each template image contains 8 images, and their order is described in language (e.g. top left to right, bottom left to right) and the VLM outputs the selected images according to our prompting as described in Figure A.5.
Figure A.5: Prompt for Fine Keyframe Detection: The figure illustrates the input image, the prompt provided to the VLM, and the output. The input image represents a visual template composed of eight frames, and the prompt requests the three best frames along with their corresponding keywords. The output displays the top three selected frames and their associated keywords.